6. Find the dy/dt for each of the following: a) y = e2t+4 b) y = te c) y = t?e21 d) y = log(77) e) y = log(7t(1 -t) 1
Answer #1

a) y = e​​​​​​2t​​​​+4

When

Y = e​​​​​​x

dy/dx = y' = e​​​​​x d(X)/dx

dy/dt = d(e​​​​​​2t+4)/dt

Y = e​​​​​​2t+4 (d(2t+4)/dt) = 2 e​​​​​​2t+4

b) y = t e​​​​​​t

dy/dt = Y = e​​​​​​t (dt/dt) + t (d(e​​​​​​t)/dt)

Y = e​​​​​​t + t e​​​​​​t

C) y = t​​​​​​2​ e​​​2t

dy/dx = Y = t​​​​2 (d(e​​​​​​2t​)/dt) + e​​​​​​2t (d(t​​​2 )/dx)

Y = t​​​​​​2 2e2t + e​​​​​​2t 2t

d) y = log (7 t​​​​​​5​​​) = log 7 + log t​​​​​​5​​​​​t5

y = log 7 + 5 log t

dy/dx = Y = (d log7/dt) + 5 ( d (log t )/dt )

Y = 0 + 5/t

Y = 5/t